Factoring
4x^2 + 19xy + 12y^2

Answer by RAY100(1637)

4x^2 +19xy +12y^2
.
sign of 3rd term is (+) therefore both factors have same sign
.
sign of second term is (+) ,,,therefore both factors are (+)
.
factors of 1st term are,, 1*4,,2*2
.
Factors of 3rd term are, 1*12,,2*6,,3*4,,
.
Seeking a combination whose factors add to 19
.
Try,,,4*4 + 1*3 = 16+3 = 19
.
(4x +3y)(1x +4y)
.